d+r=13
1.11r+0.95d=13.15

Answer by Alan3354(69443)

d + r = 13 --> r = 13-d
1.11r + 0.95d = 13.15
-----
Multiply by 100
111r + 95d = 1315
Sub for r
111*(13-d) + 95d = 1315
1443 - 111d + 95d = 1315
-16d = -128
d = 8
================
r = 5